Amber completed a 25-kilometer road race in 3 hours. She ran at the same speed and covered the same distance each hour. Between which 2 whole numbers does the distance she ran each hour lie? Draw a visual model to find the distance she ran each hour. You don’t have to do the last part but it’d be great if you did
1
Expert's answer
2019-04-24T03:52:44-0400

Amber covered each hour the same distance. So, each of three hours she covered third part from the total distance.


"S_{per hour}=\frac{S}{3}=\frac{25}{3}=8 \frac{1}{3}"

This number lies between 8 and 9.

Answer: the distance Amber ran each hour lie between 8 and 9.


To solve the task graphically we can use the following graph.




Let's place horizontally the segment (a), divided into 25 equal parts. It represents the distance ran.

At an acute angle, we construct another segment (b), which consists of three equal parts. Connect the ends of the segments with a line (green).

Parallel straight lines, splitting the segment b into three equal parts, divide the segment a and into three equal parts. Then we can calculate how many parts each of the broken segments consist of (more than 8, less than 9).